12. Predict the signs of the AG and AH for the following: a) Vaporization of water at 100°C. b) Vaporization of water above 100°C BONUS: 35 mL of 0.0025M KoCrO4 are MIXED with 75 mL of 0.000220M Pb(NOs)2. . Hint: Determine your overall and net ionic equations first!
Answer #1

(a)

Vaporisation of water at 1000C,

deltaH = + ve

deltaG = 0 SInce at this state the system is at equilibrium.

(b)

Vaporisation of water at above 1000C

deltaH = + ve

deltaG = - Ve

****

New, [Pb2+] = 0.000220 * 75 / 110 = 0.000150 M

[CrO42-] = 0.00250 * 35 / 110 = 0.000795 M

Solubility equation,

PbCrO4 (s) \rightleftharpoons Pb2+(aq.) + CrO42- (aq.)

Solubility product expression,

Qsp = [Pb2+][CrO42-]

Qsp = ( 0.000150 ) * ( 0.000795 )

Qsp = 1.19 * 10-7

Since Q sp > Ksp , precipitation occurrs.